# Training LLMs over Neurally Compressed Text

### v2

@@ Abstract
In this paper, we explore the idea of training large language models (LLMs) over highly compressed text. While standard subword tokenizers compress text by a small factor, neural text compressors can achieve much higher rates of compression. If it were possible to train LLMs directly over neurally compressed text, this would confer advantages in training and serving efficiency, as well as easier handling of long text spans. The main obstacle to this goal is that strong compression tends to produce opaque outputs that are not well-suited for learning. In particular, we find that text naïvely compressed via Arithmetic Coding is not readily learnable by LLMs. To overcome this, we propose Equal-Info Windows, a novel compression technique whereby text is segmented into blocks that each compress to the same bit length. Using this method, we demonstrate effective learning over neurally compressed text that improves with scale, and outperforms byte-level baselines by a wide margin on perplexity and inference speed benchmarks. While our method delivers worse perplexity than subword tokenizers for models trained with the same parameter count, it has the benefit of shorter sequence lengths. Shorter sequence lengths require fewer autoregressive generation steps, and reduce latency. Finally, we provide extensive analysis of the properties that contribute to learnability, and offer concrete suggestions for how to further improve the performance of high-compression tokenizers.
@@ 2.4 Arithmetic Coding
The coding component of Arithmetic Coding is nearly optimal: the output bitstream will have a length of $-\lceil\log\hat{p}(x_{0:N})\rceil+1$ bits when using infinite precision. In the finite precision setting using $\beta$ bits, an extra $O(N2^{-\beta})$ bits are added (Howard & Vitter, 1992). See (Witten et al., 1987) for an example implementation. In our experiments, we use precision $\beta=14$. The practical effect of using a finite precision implementation of Arithmetic Coding is that the model’s cumulative distribution gets quantized to integers using $\beta$ bits. This results in a minimum probability of $2^{-\beta}$ being assigned to all tokens.
@@ 2.5 Related Work
Recent work has looked at using large language models for compression, but has not to our knowledge attempted to train subsequent models over the resulting compressed output. Works like (Delétang et al., 2024) use a transformer language model as the modeling component of Arithmetic Coding, but they do not train over compressed output nor do they make modifications to the compression algorithm to facilitate learnability by downstream models. Additionally, they focus on the setting of compressing fixed-size sequences of bytes. By contrast, our models operate over input sequences of fixed token length. This allows for models with higher compression rates to leverage longer contexts, as more bytes are included in the input.
(Valmeekam et al., 2023) proposes changes to Arithmetic Coding to make it more amenable to use with LLMs—namely, they rank sort the logits from the model before creating text intervals, $I_{i}(x_{0:N})$. This could help alleviate issues stemming from errors in M2’s simulation of M1. However, they do not train models on top of their compressed output.
@@
Other recent work focuses on using the “modeling” component from well-known compressors to do other tasks. (Jiang et al., 2022) uses the model from GZip to perform text classification. (Vilnis et al., 2023) uses the Arithmetic Decoding algorithm with an LLM as the model to do diverse parallel sampling from that LLM. One could imagine that the “model” of our compressors (M1) is a teacher for M2, but unlike these other applications, the M1 values are not used outside of compression.
(Külekci, 2011) also explores learning over compressed text, but with several key differences. First, they use n-gram language models (Shannon, 1948) while we use LLMs. Second, their model is conditioned on compressed bitstreams but produces a distribution over the raw, uncompressed, bytes while our M2 models predict directly in the compressed space. Additionally, they only consider static Huffman coding (Huffman, 1952) as the algorithm to compress model inputs. While this avoids the context sensitivity issues we outline in Section 2.2, it results in a far worse compression rate compared to the adaptive compression methods we use. One important distinction is that their equal-information windows are overlapping, and used as a sliding window to provide context to their n-gram language model. By contrast our equal-information windows are non-overlapping, and used to segment text into a series of equal-length bitstrings that can be interpreted independently by M2, and whose boundaries are easily identifiable, as they map to a fixed number of M2 tokens.
Concurrently, (Goldman et al., 2024) explores how the compression performance of a tokenizer correlates with downstream model performance. They find that tokenizers that compress better perform better, which generally aligns with our findings, particularly in the large vocabulary setting, see Fig. 6. However, we find that using the strongest compressors is detrimental to learnability, as seen in the AC line in Fig. 3. These conflicting results likely stem from differences in tokenization strategy. Their work is restricted to BPE-based compressors while we explore stronger compressors built on LLMs and Arithmetic Coding. The qualitative differences between these classes of tokenizers are explored more in Section 6.1.
@@ 3.3 Compression Methods
The compressor model is run over contiguous text sequences of $10{,}240$ bytes. The generated logits are used as the model distribution for Arithmetic Coding. We use the Range Encoding (a finite-precision implementation of Arithmetic Coding) implementation from TensorFlow Compression (Ballé et al., 2024) with a precision of $14$. The range encoding implementation uses integers with precision+ $+\,2$ bits. This is enough to encode $16$-bit float logits, so should not cause numerical issues as our models are trained using bfloat16. While the compressor model is only trained on sequences of length $1{,}024$, it uses relative position embeddings in its attention layers. Thus, it can be applied to longer sequences. Some works observe decreased performance as inputs are scaled to lengths beyond those seen in training (Varis & Bojar, 2021, Press et al., 2022), but we find that compression performance is similar in the two settings. Compressing sequences of length $1{,}024$ yields a compression ratio of $5.46$ while compressing sequences of length $10{,}240$ yields a ratio of $5.49$. This suggests the performance drop from long sequences has minimal effect on compression, or that the increased contextual information makes up this difference.
@@
To weaken the compressor model, we replace the context-sensitive LM model with a static byte unigram model—that is, the model’s distribution is the same for all byte tokens in the input, i.e., $\hat{p}(x_{i}|x_{0},\dots,x_{i-1})=\hat{p}(x_{i})$. This distribution is estimated using the byte unigram statistics from the C4 training data.
@@ 3.6 Baselines
Bytes: These baselines train directly over UTF-8 bytes, using the byte tokenizer from ByT5 (Xue et al., 2022). The models see $26.2$ billion bytes total (see Table 12).
SentencePiece: These baselines train on text tokenized with the SentencePiece vocabulary of $32{,}000$ tokens from T5 (Raffel et al., 2020). The models see $112$ billion bytes total (see Table 12).
@@ 3.8 Evaluation
We also compare models on how much computation (FLOPs) is required to perform inference over a given length of raw text (bytes). More specifically, we calculate M2’s expected FLOPs/byte by scaling FLOPs/token—approximated by $2\,\times\,\texttt{params}$ (excluding embedding parameters) following (Kaplan et al., 2020)—by the token-level compression rate (as tokens/byte). For methods using an M1 model during compression, the FLOPs/byte cost of M1 is added. [^10] While there is a computational cost to running GZip over the input text, we ignore it as it is insubstantial compared to the cost of running M2 model inference. For more details on the evaluation metrics see Section D.4.
@@ 4 Results
Figure 3: Models trained over compressed text are compared against baseline models in terms of bits/byte ($\downarrow$) and inference FLOPs/bytes ($\downarrow$). The ArithmeticCoding and StaticAC settings are essentially unlearnable, with models failing to outperform naïve baselines (dashed lines) that assign equal probability to all tokens. EqualInfoAC and GZip outperform naïve baselines and show improvement with scale. EqualInfoAC is the strongest of the compression-based methods, with EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforming the Bytes baseline at all sizes. While SentencePiece performs the best, the gap between EqualInfoAC and SentencePiece narrows with scale. See Section D.9 for the exact values used in this and other graphs.
@@
Our SentencePiece baseline outperforms all other methods, including our Bytes baseline, across all model sizes. On the surface, this result seems to run counter to the recent findings of (Delétang et al., 2024), where their byte-level models outperformed subword (BPE) models at medium and large scales. The discrepancy is due to prioritizing different metrics. They report the model’s bit compression rate on fixed-length ($2{,}048$ byte) sequences. While this is one type of “fair” comparison, it disadvantages subword models, as they are trained to model dependencies longer than $2{,}048$ bytes (but never evaluated on this ability), and are allotted fewer inference FLOPs to process the same text, as compared to the byte-level models. Additionally, bit compression ratio penalizes subword models for having larger vocabulary sizes. By contrast, our evaluation tests what perplexity models achieve on sequences of the same length they were trained on, and compares models at matching FLOPs/byte cost. This aligns with our end goal, which is to train an LLM that achieves the best perplexity at whatever sequence length it can handle, given a fixed budget for training and inference.
@@
However, EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforms the SentencePiece baseline in terms of tokens/byte. Models using EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ take fewer autoregressive steps to generate the same text than models using SentencePiece encoding. This has the potential to reduce generation latency, at the cost of reduced compute efficiency. This is a tradeoff that is often worth making in production. For instance, speculative decoding (Leviathan et al., 2023b) is a popular approach that performs redundant computation in order to potentially accelerate auto-regressive steps.
@@
It is apparent from Fig. 3 that if FLOPs/byte were held constant, SentencePiece would achieve slightly better bits/byte than EqualInfoAC. However there is another axis along which EqualInfoAC may still be preferred. Setting aside inference FLOPs, all our SentencePiece models require $23$% longer sequences to encode the same text when compared to our best EqualInfoAC setting ($b\mathord{=}16$, $v\mathord{=}65$k). This means that regardless of FLOPs used, the SentencePiece models will take more decoder steps at inference time. It is up to the practitioner whether it is “worth it” to trade off some bits/byte performance in order to achieve shorter sequences. In many serving scenarios, decoder steps are a practical bottleneck for determining system latency, and there are cases where one may be willing to incur extra inference cost to reduce latency (e.g., speculative decoding (Leviathan et al., 2023a)). To this end, it may be advantageous to scale up an EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ model to recover bits/byte performance while retaining the reduced latency. This can be seen visually in Fig. 4.
@@
Given the recent findings of (Schaeffer et al., 2023), we anticipate that continuing to scale models beyond $2$ billion parameters is unlikely to deliver an “emergent” ability to learn over AC-compressed text, since the bits/byte metric we use is smooth.
@@
When scaling models, (Hoffmann et al., 2022) recommend that training tokens should be scaled linearly with model size. However, in our experiments above, all models see the same number of tokens, regardless of model size. Consequently, our largest models may be somewhat “undertrained”. [^12] The undertraining of our $2$b models is also visible in their validation loss curves, which still have a significant decreasing slope at $200{,}000$ steps, showing the models have not yet converged. To test whether following the “scaling laws” recommendation influences our results, we reevaluate our models at earlier checkpoints selected to maintain a constant ratio of training data to model size. We find that all core trends are unchanged in this setting. See Section D.3 for details.
@@ 5.2 Transformers struggle to learn Arithmetic Coding
EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$ Decompression $95.30$ $0.141\pm 0.002$
@@
AC$[v\mathord{=}256]$ Compression $1.7$ $2.489$
To directly diagnose the ability to track Arithmetic Coding, we format AC compression and decompression as sequence-to-sequence tasks. The input provides the model with the true text, so we expect a model that is able to learn Arithmetic Coding should achieve an accuracy of $100$. We compress sequences of $1{,}024$ bytes using M1 and Arithmetic Coding. [^13] We use shorter raw text sequences to keep the final sequence length of inputs + targets manageable. We concatenate the bytes and AC output tokens to create the compression task. For the decompression task, we simply flip the order—AC output tokens first and then bytes. The target token IDs (bytes or tokens) are shifted by the input vocabulary size, ensuring that they have distinct values. We use a decoder-only transformer as our model with a causal attention mask, i.e., even during the input sequence, future tokens are hidden from the model. We train models with $113$m parameters. Loss, gradients, and evaluation metrics are only computed on the target tokens.

### v3

@@ Abstract
In this paper, we explore the idea of training large language models (LLMs) over highly compressed text. While standard subword tokenizers compress text by a small factor, neural text compressors can achieve much higher rates of compression. If it were possible to train LLMs directly over neurally compressed text, this would confer advantages in training and serving efficiency, as well as easier handling of long text spans. The main obstacle to this goal is that strong compression tends to produce opaque outputs that are not well-suited for learning. In particular, we find that text naïvely compressed via Arithmetic Coding is not readily learnable by LLMs. To overcome this, we propose Equal-Info Windows, a novel compression technique whereby text is segmented into blocks that each compress to the same bit length. Using this method, we demonstrate effective learning over neurally compressed text that improves with scale, and outperforms byte-level baselines by a wide margin on perplexity and inference speed benchmarks. While our method delivers worse perplexity than subword tokenizers for models trained with the same parameter count, it has the benefit of shorter sequence lengths. Shorter sequence lengths require fewer autoregressive generation steps, often reducing latency. Finally, we provide extensive analysis of the properties that contribute to learnability, and offer concrete suggestions for how to further improve the performance of high-compression tokenizers.
@@ 2.4 Arithmetic Coding
The coding component of Arithmetic Coding is nearly optimal: the output bitstream will have a length of $-\lceil\log\hat{p}(x_{0:N})\rceil+1$ bits when using infinite precision. In the finite precision setting using $\beta$ bits, an extra $O(N2^{-\beta})$ bits are added (Howard & Vitter, 1992). See Witten et al. (1987) for an example implementation. In our experiments, we use precision $\beta=14$. The practical effect of using a finite precision implementation of Arithmetic Coding is that the model’s cumulative distribution gets quantized to integers using $\beta$ bits. This results in a minimum probability of $2^{-\beta}$ being assigned to all tokens.
@@ 2.5 Related Work
Recent work has looked at using large language models for compression, but has not to our knowledge attempted to train subsequent models over the resulting compressed output. Works like Delétang et al. (2024) use a transformer language model as the modeling component of Arithmetic Coding, but they do not train over compressed output nor do they make modifications to the compression algorithm to facilitate learnability by downstream models. Additionally, they focus on the setting of compressing fixed-size sequences of bytes. By contrast, our models operate over input sequences of fixed token length. This allows for models with higher compression rates to leverage longer contexts, as more bytes are included in the input.
Valmeekam et al. (2023) proposes changes to Arithmetic Coding to make it more amenable to use with LLMs—namely, they rank sort the logits from the model before creating text intervals, $I_{i}(x_{0:N})$. This could help alleviate issues stemming from errors in M2’s simulation of M1. However, they do not train models on top of their compressed output.
@@
Other recent work focuses on using the “modeling” component from well-known compressors to do other tasks. Jiang et al. (2022) uses the model from GZip to perform text classification. Vilnis et al. (2023) uses the Arithmetic Decoding algorithm with an LLM as the model to do diverse parallel sampling from that LLM. One could imagine that the “model” of our compressors (M1) is a teacher for M2, but unlike these other applications, the M1 values are not used outside of compression.
Külekci (2011) also explores learning over compressed text, but with several key differences. First, they use n-gram language models (Shannon, 1948) while we use LLMs. Second, their model is conditioned on compressed bitstreams but produces a distribution over the raw, uncompressed, bytes while our M2 models predict directly in the compressed space. Additionally, they only consider static Huffman coding (Huffman, 1952) as the algorithm to compress model inputs. While this avoids the context sensitivity issues we outline in Section 2.2, it results in a far worse compression rate compared to the adaptive compression methods we use. One important distinction is that their equal-information windows are overlapping, and used as a sliding window to provide context to their n-gram language model. By contrast our equal-information windows are non-overlapping, and used to segment text into a series of equal-length bitstrings that can be interpreted independently by M2, and whose boundaries are easily identifiable, as they map to a fixed number of M2 tokens.
Several previous works explore how a tokenizer’s compression rate correlates with downstream model performance. Goldman et al. (2024) finds that tokenizers that compress better perform better, which generally aligns with our findings, particularly in the large vocabulary setting, see Fig. 6. However, we find that using the strongest compressors is detrimental to learnability, as seen by the AC line in Fig. 3.
By contrast, Schmidt et al. (2024) and Dagan et al. (2024) observe a weak trend in the opposite direction—better downstream performance from tokenizers that compress less. This discrepancy may be due to the fact that Goldman et al. (2024) artificially weakens compression by undertraining the tokenizer, whereas Dagan et al. (2024) compares existing popular tokenizers, and Schmidt et al. (2024) compares different subword algorithms.
The divergences between these results and ours likely stem from major differences in tokenization strategy. These three works are restricted to subword compressors, whereas we explore stronger compressors built on LLMs and Arithmetic Coding. Among other differences, subword tokenizers output tokens in a near-Zipfian distribution (Zipf, 1935), whereas our compressors output a near-uniform distribution, see Table 3. The qualitative differences between these classes of tokenizer are explored more in Section 6.1.
Rajaraman et al. (2024) and Makkuva et al. (2024) find that under some data generation paradigms, transformers fail to model the contextual generation process and instead learn the underlying stationary distribution of the process. This is similar to the failure case we observe where M2 models only output a uniform distribution after training on compressed text. However, their synthetic data setting is quite different from our compressed data setting.
@@ 3.3 Compression Methods
The compressor model is run over contiguous text sequences of $10{,}240$ bytes. The generated logits are used as the model distribution for Arithmetic Coding. We use the Range Encoding (a finite-precision implementation of Arithmetic Coding) implementation from TensorFlow Compression (Ballé et al., 2024) with a precision of $14$. The range encoding implementation uses integers with precision+ $+\,2$ bits. This is commonly used when encoding $16$-bit float logits, so we do not expect it to cause numerical issues as our models are trained using bfloat16. While the compressor model is only trained on sequences of length $1{,}024$, it uses relative position embeddings in its attention layers. Thus, it can be applied to longer sequences. Some works observe decreased performance as inputs are scaled to lengths beyond those seen in training (Varis & Bojar, 2021, Press et al., 2022), but we find that compression performance is similar in the two settings. Compressing sequences of length $1{,}024$ yields a compression ratio of $5.46$ while compressing sequences of length $10{,}240$ yields a ratio of $5.49$. This suggests the performance drop from long sequences has minimal effect on compression, or that the increased contextual information makes up this difference.
@@
To weaken the compressor model, we replace the context-sensitive LM model with a static byte unigram model—that is, the model’s distribution over bytes is the same for each token in the input, i.e., $\hat{p}(x_{i}|x_{0},\dots,x_{i-1})=\hat{p}(x_{i})$. This distribution is estimated using the byte unigram statistics from the C4 training data.
@@ 3.4 Tokenization of Compressed Text
Following Rajaraman et al. (2024), our tokenization scheme can be described as the tuple $\mathcal{T}=(\mathrm{Dict},\mathrm{DS},\mathrm{enc}(\cdot),\mathrm{dec}(\cdot))$. $\mathrm{Dict}$ is the space of tokens created by segmenting the compressed bitstream—in our case, a vocabulary of $256$ or $65$k. $\mathrm{DS}$ is the entire M1 model. The functions $\mathrm{enc}(\cdot)$ and $\mathrm{dec}(\cdot)$ perform encoding (compression) and decoding (decompression). In our case, these functions vary with (i) the M1 model, (ii) the compression algorithm (AC, EqualInfoAC, etc.), and (iii) the bitstream segmentation strategy.
@@ 3.6 Baselines
Bytes: These baselines train directly over UTF-8 bytes, using the byte tokenizer from ByT5 (Xue et al., 2022), which simply encodes the input string as UTF-8 using Python s.encode("utf-8"). The models see $26.2$ billion bytes total (see Table 12).
SentencePiece: These baselines train on text tokenized using the T5 (Raffel et al., 2020) vocabulary, which has $32{,}000$ tokens, and was trained using the Unigram algorithm (Kudo, 2018) as implemented by the SentencePiece library (Kudo & Richardson, 2018). Unigram is one of several popular subword tokenization algorithms, with others including BPE (Sennrich et al., 2016) and WordPiece (Schuster & Nakajima, 2012, Wu et al., 2016), and each of these is implemented by multiple libraries, sometimes with minor differences.
We opt to use the SentencePiece Unigram tokenizer as our subword baseline for a few reasons. First, the performance of these competing subword tokenization algorithms is known to be very similar, with several works finding no statistically significant difference between them in many settings (Kudo, 2018, Ali et al., 2024, Schmidt et al., 2024). Second, where differences are reported (marginal or otherwise), Unigram has been found to be preferred, at least for English-only models (Kudo, 2018, Bostrom & Durrett, 2020, Ali et al., 2024, Schmidt et al., 2024). Finally, previous work indicates that the SentencePiece implementation of Unigram outperforms the HuggingFace implementation (Ali et al., 2024, Schmidt et al., 2024).
Our baseline models trained over SentencePiece tokens see $112$ billion bytes total (see Table 12).
@@ 3.8 Evaluation
We also compare models on how much computation (FLOPs) is required to perform inference over a given length of raw text (bytes). More specifically, we calculate M2’s expected FLOPs/byte by scaling FLOPs/token—approximated by $2\,\times\,\texttt{params}$ (excluding embedding parameters) following Kaplan et al. (2020)—by the token-level compression rate (as tokens/byte). For methods using an M1 model during compression, the FLOPs/byte cost of M1 is added. [^10] While there is a computational cost to running GZip over the input text, we ignore it as it is insubstantial compared to the cost of running M2 model inference. For more details on the evaluation metrics see Section D.4.
@@ 4 Results
Figure 3: Models trained over compressed text are compared against baseline models in terms of bits/byte ($\downarrow$) and inference FLOPs/byte ($\downarrow$). The ArithmeticCoding and StaticAC settings are essentially unlearnable, with models failing to outperform naïve baselines (dashed lines) that assign equal probability to all tokens. EqualInfoAC and GZip outperform naïve baselines and show improvement with scale. EqualInfoAC is the strongest of the compression-based methods, with EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforming the Bytes baseline at all sizes. While SentencePiece performs the best, the gap between EqualInfoAC and SentencePiece narrows with scale. See Section D.9 for the exact values used in this and other graphs.
@@
On the surface, M2’s inability to learn seems to run counter to Yun et al. (2020), which finds that transformers are universal approximators of sequence-to-sequence functions. However, the fact that transformers can express a function does not yet imply the ability to learn that function via stochastic gradient descent from random initialization. Additionally, for the purpose of training LLMs, our interest is in training a model that generalizes well to held-out validation data, as opposed to one that overfits its training data. In practice, we find our M2 training procedure is sufficient to overfit to small subsets of our AC-compressed training data, when repeated over multiple epochs. However, this does not satisfy our goal of finding a compression scheme suitable for LLM training.
@@
Our SentencePiece baseline outperforms all other methods, including our Bytes baseline, across all model sizes. On the surface, this result seems to run counter to the recent findings of Delétang et al. (2024), where their byte-level models outperformed subword (BPE) models at medium and large scales. The discrepancy is due to prioritizing different metrics. They report the model’s bit compression rate on fixed-length ($2{,}048$ byte) sequences. While this is one type of “fair” comparison, it disadvantages subword models, as they are trained to model dependencies longer than $2{,}048$ bytes (but never evaluated on this ability), and are allotted fewer inference FLOPs to process the same text, as compared to the byte-level models. Additionally, bit compression ratio penalizes subword models for having larger vocabulary sizes. By contrast, our evaluation tests what perplexity models achieve on sequences of the same length they were trained on, and compares models at matching FLOPs/byte cost. This aligns with our end goal, which is to train an LLM that achieves the best perplexity at whatever sequence length it can handle, given a fixed budget for training and inference.
@@
However, EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforms the SentencePiece baseline in terms of tokens/byte. Models using EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ take fewer autoregressive steps to generate the same text than models using SentencePiece encoding. This has the potential to reduce generation latency, at the cost of reduced compute efficiency. This is a tradeoff that is often worth making in production. For instance, speculative decoding (Leviathan et al., 2023) is a popular approach that performs redundant computation in order to potentially accelerate auto-regressive steps.
@@
It is apparent from Fig. 3 that if FLOPs/byte were held constant, SentencePiece would achieve slightly better bits/byte than EqualInfoAC. However there is another axis along which EqualInfoAC may still be preferred. Setting aside inference FLOPs, on average SentencePiece tokenization requires $23$% longer sequences to encode the same text when compared to our best EqualInfoAC setting ($b\mathord{=}16$, $v\mathord{=}65$k). This means that, regardless of FLOPs used, the SentencePiece models will take more decoder steps at inference time. It is up to the practitioner whether it is “worth it” to trade off some bits/byte performance in order to achieve shorter sequences. In many serving scenarios, decoder steps are a practical bottleneck for determining system latency, as other aspects of model scale, such as width, can be mitigated with enough parallelism. There are cases where one may be willing to incur even more (parallelizable) inference costs to reduce latency, as in “speculative decoding” (Leviathan et al., 2023). To this end, it may be advantageous to use more compute resources to scale up an EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ model (recovering bits/byte performance) while retaining the reduced latency due to the shorter sequence length. This can be seen visually in Fig. 4.
@@
Given the recent findings of Schaeffer et al. (2023), we anticipate that continuing to scale models beyond $2$ billion parameters is unlikely to deliver an “emergent” ability to learn over AC-compressed text, since the bits/byte metric we use is smooth.
@@
When scaling models, Hoffmann et al. (2022) recommend that training tokens should be scaled linearly with model size. However, in our experiments above, all models see the same number of tokens, regardless of model size. Consequently, our largest models may be somewhat “undertrained”. [^12] The undertraining of our $2$b models is also visible in their validation loss curves, which still have a significant decreasing slope at $200{,}000$ steps, showing the models have not yet converged. To test whether following the “scaling laws” recommendation influences our results, we reevaluate our models at earlier checkpoints selected to maintain a constant ratio of training data to model size. We find that all core trends are unchanged in this setting. See Section D.3 for details.
@@ 5.2 Transformers Struggle to Learn Arithmetic Coding
Figure 7: Arithmetic compression and decompression cast as sequence-to-sequence tasks. To account for the strong performance that is possible by modeling only the output bytes in the decompression task, we also train a Byte LM on just the targets.
@@
EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$ Decompress $95.30$ $0.141\pm 0.002$
@@
AC$[v\mathord{=}256]$ Compress $1.7$ $2.489$
To directly diagnose the ability to track Arithmetic Coding, we format AC compression and decompression as sequence-to-sequence tasks, as shown in Fig. 7. The input provides the model with the true text, so we expect a model that is able to learn Arithmetic Coding should achieve an accuracy of $100$. We compress sequences of $1{,}024$ bytes using M1 and Arithmetic Coding. [^13] We use shorter raw text sequences to keep the final sequence length of inputs + targets manageable. We concatenate the bytes and AC output tokens to create the compression task. For the decompression task, we simply flip the order—AC output tokens first and then bytes. The target token IDs (bytes or tokens) are shifted by the input vocabulary size, ensuring that they have distinct values. We use a decoder-only transformer as our model with a causal attention mask, i.e., even during the input sequence, future tokens are hidden from the model. We train models with $113$m parameters. Loss, gradients, and evaluation metrics are only computed on the target tokens.